OpenStudy (anonymous):

what are the possible roots for 3x^2-3x-5

OpenStudy (anonymous):

hey is this algebra 2

OpenStudy (anonymous):

x = 1/6 (3+sqrt(69)) x = 1/6 (3-sqrt(69))